Find the area of the surface that is generated by revolving the portion of the curve y=x^2 between x=0 and x=1 about the y-axis.


1
Expert's answer
2021-11-18T17:54:43-0500

Solution:

The curve y = "\ x^{2}" from x = 0 and x = 1 about the y axis


We know that x = 0 to x = 1 , which is analogous from y = 0 to y = 1.

Area of the surface is

"A = 2\times \pi \int_{a}^{b} g(y)\sqrt{1+(g'(y)^{2})}dy"


"g(y) = \sqrt{y}"

"g'(y) = \frac {1}{(2\times\sqrt{y})}"

"A = 2\times \pi \int_{0}^{1} \sqrt {y} \sqrt{1+(\frac{1}{(2\times\sqrt{y})})^{2}}dy"

"\Rightarrow A = 2\times \pi \int_{0}^{1} \sqrt{y(1+\frac{1}{4y})}dy"

"\Rightarrow A = 2\times \pi \int_{0}^{1} \sqrt{y+(\frac{1}{4})}dy"

Let,

"u = y + \frac {1}{4}"


"\Rightarrow du = dy"


"y = 0 , u = \frac {1} {4}"


"y = 1 , u = \\frac {5} {4}\n\n\u200b"


"\Rightarrow A = 2\times \pi \int_{\frac{1}{4}}^{\frac{5}{4}} u^ { \frac {1} {2} } du"


"\Rightarrow A= 2 \times \pi [\frac{2}{3} (u^\frac {3} {2}) ]|_{\frac{1}{4}} ^{ \frac{5}{4}}"


"\Rightarrow A= \frac{4 \times \pi}{3} [(\frac{5}{4})^\frac {3} {2} - (\frac{1}{4})^\frac {3} {2}]"


"\Rightarrow A= 5.33 sq.unit"